• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 08:23:17 on May 10, 2012

    Entry number 383555

    keyword=Owl Shift Summary

    Owl Shift Summary
    
    SL: Melissa Cummings
    TO: Yuxiang Zhao
    3rd: Aji Daniel
    
    
    00:00 Pengjia is starting harp scans
    00:18 Magnet quench (See halog 383540).
    00:20 Called for escorted access.  James, Dustin and Josh  went down to
    the hall to look at the magnet.
    02:30 Work on target is finished, magnet is cooling down
    03:30 Refill of liquid helium is complete and magnet has cooled down. 
    James came in to ramp the magnet back up.  
    04:15 Target magnet is ramped to ~100A.  James went down to the hall to
    tune the NMR while the magnet finishes ramping up.
    05:00 NMR tune is done. Taking baselines in preparation for TEs
    05:20 Starting TEs
    08:00 End of shift - still TE'ing